OpenStudy (anonymous):

what is the value of cosine 14pi/3

OpenStudy (waleed_imtiaz):

Write it as Cos 840=cos (9(90)+30) =sin 30=1/2...... Cosine changes to sine because there's an odd multiple of pi/2 and the answer is in positive sign because 840 degrees lies in 2nd quadrant and sine is positive in 2nd quadrant..... U got it ?